OpenStudy (anonymous):

how do I find the area of 2 1/5 base yards, height 5/9yards; ok so trying to work this, is this correct??? (2 1/5*5/9)/2; (11/5* 5/9)/2; (55/45)(1/2) 55/90????But after that stuck

OpenStudy (anonymous):

ok so trying to work this, is this correct??? (2 1/5*5/9)/2; (11/5* 5/9)/2; (55/45)(1/2) 55/90????But after that stuck

OpenStudy (anonymous):

the answer is \[\frac{ 11 }{ 9 }\] because area=length*width so you would do \[\frac{ 11 }{ 5 }\times \frac{ 5 }{ 9 }=\frac{ 55 }{ 45 }=\frac{ 11 }{ 9}\]